In a Beams ultracentrifuge, the rotor is suspended magnetically in a vacuum. Since there is no mechanical connection to the rFor a point at the end of the rotor, find the magnitude of the tangential acceleration component. m/s2For a point at the end of the rotor, find the maximum radial acceleration component. m/s2
Answer #1

Initial angular speed, ω = 5.0 x 10^5 rad/s

Radius of the centrifuge = length of the rod = r = 38.2 cm = 0.382 m

(a) Initial speed, v = r*ω

= 0.382 m x 5.0 x 10^5 rad/s

= 1.91 x 10^5 m/s (Answer)

(b) Tangential acceleration component, a(tan) = r*α

= 0.382 m*(-0.40 rad/s^2) = -0.1528 m/s^2 (Answer)

(c) Maximum radial acceleration component, a(rad) = r*ω^2

= 0.382 m*(5.0x10^5 rad/s)^2

= 9.55 x 10^10 m/s^2 (Answer)